Extremely confused daughter tried to help and she is stumped this is from an algebra practice test word problem is:
A merchant has coffee worth $10 a pound that she wants to mix with 40 pounds of coffee worth $80 a pound to get a mixture that can be sold at $30 a pound. How many pounds of the $10 coffee should be use. Can you explain and show how you came up with answer please. Thank you for any help you can give us.

Found 2 solutions by mananth, shree840:
Answer by mananth(16949)

------ Price ---------------- quantity
Coffee type I 80 ---------------- 40
Coffee type II 10 ---------------- x
Total 30 ---------------- 40 + x

80*40+10*x=30(40+x)
3200+10x=1200+30x
10x-30x=1200-3200
-20x=-2000
/-20 \
x= 100 pounds 10% Coffee type II

Answer by shree840(260)

let us use x pounds of coffee total cost 10x cost of 40 lbs at $80=3200
total pounds=40+x total cost = 30(40+x)
10x+3200=30(40+x)=1200+30x
2000=30x-10x=20x
100=x
so we mix 100lbs at $10 with 40lbs at $80
total cost 3200+1000=4200
total lbs =100+40=140
cost =$30 Ans